# Galaxy evolution in the post-merger regime. III - The triggering of active galactic nuclei peaks immediately after coalescence

### v1

@@ Abstract
Galaxy mergers have been shown to trigger active galactic nuclei (AGN) in the nearby universe, but the timescale over which this process happens remains unconstrained. The Multi-Model Merger Identifier (mummi) machine vision pipeline has been demonstrated to provide reliable predictions of time post-merger ($T_{PM}$) for galaxies selected from the Ultraviolet Near Infrared and Optical Northern Survey (UNIONS) up to $T_{PM}$=1.76 Gyr after coalescence. By combining the post-mergers identified in UNIONS with pre-coalescence galaxy pairs, we can, for the first time, study the triggering of AGN throughout the merger sequence. AGN are identified using a range of complementary metrics: mid-IR colours, narrow emission lines and broad emission lines, which can be combined to provide insight into the demographics of dust and luminosity of the AGN population. Our main results are: 1) Regardless of the metric used, we find that the peak AGN excess (compared with a matched control sample) occurs immediately after coalescence, at $0.160<T_{PM}<$ 0.16 Gyr. 2) The excess of AGN is observed until long after coalescence; both the mid-IR selected AGN and broad line AGN are more common than in the control sample even in the longest time bin of our sample ($0.96<T_{PM}$ < 1.76 Gyr). 3) The AGN excess is larger for more luminous and bolometrically dominant AGN, and we find that AGN in post-mergers are generally more luminious than secularly triggered events. 4) A deficit of broad line AGN in the pre-merger phase, that evolves into an excess in post-mergers is consistent with evolution of the covering fraction of nuclear obscuring material. Before coalescence, tidally triggered inflows increase the covering fraction of nuclear dust; in the post-merger regime feedback from the AGN clears (at least some of) this material. 5) The statistical peak in the triggering of starbursts occurs contemporaneously with AGN, within 0.16 Gyr of coalescence.
@@ 1 Introduction
For many years, the link between galaxy mergers and the triggering of nuclear accretion was considered highly contentious. Indeed, it had become an almost customary component of any introduction on the subject to point out the slew of papers that both found support for (e.g. Alonso et al. 2007; Woods & Geller 2007; Koss et al. 2010; Ellison et al. 2011; Liu et al. 2012; Sabater et al. 2013; Khabiboulline et al. 2014; Satyapal et al. 2014) and against (e.g. Ellison et al. 2008; Li et al. 2008; Gabor et al. 2009; Darg et al. 2010; Cisternas et al. 2011; Scott & Kaviraj 2014) a link between galaxy mergers and a statistical excess of active galactic nuclei (AGN). Since these pioneering works of over a decade ago, we have learned many lessons, including the importance of large statistical datasets, rigourously matched control samples and multi-wavelength AGN selection. Thus, although the situation at $z>0.5$ is still debated (e.g. Kocevski et al. 2012; Shah et al. 2020; Silva et al. 2021; Lambrides et al. 2021; Dougherty et al. 2024; La Marca et al. 2024), modern studies of low redshift galaxies ($z<0.5$) now generally agree that mergers are capable of triggering AGN (e.g. Weston et al. 2017; Goulding et al. 2018; Ellison et al. 2019; Gao et al. 2020; Duplancic et al. 2021; Bickley et al. 2023, 2024a; Li et al. 2023; La Marca et al. 2024).
@@
Despite the broad consensus view that low $z$ mergers can trigger AGN, many details of the merger-AGN connection remain unclear. For example, whether or not merger triggered AGN are more luminous (Koss et al. 2012; Satyapal et al. 2014; Marian et al. 2020; Bickley et al. 2023; Villforth et al. 2017; Villforth 2023; La Marca et al. 2024) or more obscured (Kocevski et al. 2015; Ricci et al. 2017; Satyapal et al. 2017; Koss et al. 2018; Li et al. 2020; Barrows et al. 2023) than secularly triggered AGN, and whether or not feedback is capable of excavating nuclear gas and dust in a so-called blowout phase (e.g. Hou et al. 2020, 2023; Ricci et al. 2021; Bickley et al. 2024a). Definitive answers to these questions remain elusive, with the further complexity that factors such as galaxy mass, star formation rate or the mass ratio of the merger may influence the outcome (e.g. Stemo et al. 2021; Barrows et al. 2023). What seems clear, however, is that the details of the sample selection, both in terms of galaxy (e.g. stellar mass, redshift, orbital geometry and merger mass ratio) and AGN properties (e.g. luminosity and obscuration) all affect the extent to which mergers and AGN are connected, as also found in simulations (e.g. Capelo et al. 2015; Blecha et al. 2018; Steinborn et al. 2018; McAlpine et al. 2020).
@@
In addition to the many galaxy-centric factors that impact the triggering of AGN during galaxy interactions, another important variable that is expected to influence experimental outcomes is time along the merger sequence. On the one hand, the time evolution of galaxy mergers is a natural axis to investigate with simulations and several works have reported a consensus view that elevated accretion can begin in the encounter phase, but peaks around the time of coalescence (Capelo et al. 2015; McAlpine et al. 2020; Byrne-Mamahit et al. 2023). Most recently, by tracking black hole accretion rates throughout the merger sequence in a large sample drawn from the Illustris-TNG simulation, Byrne-Mamahit et al. (2024) found that the statistical elevation of merger triggered nuclear accretion is long lived, persisting for $\sim$ 2 Gyr post-coalescence.
@@
On the other hand, studying merger properties as a function of time is extremely challenging for experimental works. The popular proxy for time in the pre-coalescence regime is the projected separation ($r_{p}$) between pairs of galaxies, which is statistically well motivated by simulations (e.g. Patton et al. 2024). Many observational works have found the greatest enhancement in AGN activity at small pair separations (e.g. Ellison et al. 2011; Khabiboulline et al 2014; Hou et al. 2020; Stemo et al. 2021), motivating studies to push into the post-coalescence regime. Indeed, studies of post-mergers have found that the highest rates of AGN activity occur after coalescence (Ellison et al. 2013; Satyapal et al. 2014; Bickley et al., 2023, 2024a), in agreement with simulations. However, what remains completely unconstrained in the observations is how nuclear activity evolves with time after coalescence, both in terms of the epoch at which it peaks, and for how long statistical enhancements persist. The application of artificial intelligence to not only identify mergers (Pearson et al. 2019, 2022; Ferreira et al. 2020; Bickley et al. 2021; Walmsley et al. 2022; Omori et el. 2023) but also predict the time since coalescence of a given post-merger remnant, is just now becoming feasible (Koppula et al. 2021; Pearson et al. 2024).
@@
In order to study the effects of galaxy interactions as a function of time post-merger ($T_{PM}$) we have recently developed the MUlti Model Merger Identifier (mummi), a machine vision based framework that can both find mergers and predict $T_{PM}$ (Ferreira et al. 2024a, 2025). In our first two papers in the science series, we have used mummi to investigate the post-coalescence time evolution of star formation triggering (Ferreira et al. 2024b) and quenching (Ellison et al. 2024). In the work presented here we turn our attention to AGN, in order to assess both the epoch of peak nuclear accretion as well as the duration over which the merger influences the heightened activity.
@@ 2.2 The galaxy pair sample
Our pairs sample is constructed following the methods laid out in Patton et al. (2016), except that we use our new photometrically derived stellar masses that include galaxies with a possible broad line component. The pairs sample is compiled by identifying the closest companion of each galaxy in the SDSS DR7 whose stellar mass is within a factor of 10, whose velocity separation $\Delta v<300$ km s-1 and whose projected separation is $r_{p}<100$ kpc. Furthermore, in order to have a sample that is consistent with the post-mergers (see next sub-section) we require that each galaxy in the pair have a stellar mass of log ($M_{\star}/M_{\odot})\geq 10$. There are 21,235 such galaxies that fulfill all of these criteria and that represent our pairs sample [^2] Strictly, this is a sample of galaxies with close companions, rather than a true sample of pairs, because the closest companions identified do not always find unique pairings. I.e. Galaxy A may have galaxy B as its closest companion, but galaxy B’s closest neighbour may be galaxy C. This explains why we have an odd number of galaxies in the ‘pairs’ sample. Nonetheless, we use the nomenclature ‘pairs’ as it is widely adopted in the literature..
@@ 2.3 The post-merger sample
The post-merger sample is identified by applying the mummi neural network framework to the Ultraviolet Near Infrared and Optical Northern Survey (UNIONS). Since mummi has been extensively described in Ferreira et al. (2024a, 2025) we give only the briefest of details here.
mummi is a machine vision based merger classifier that uses an ensemble of 10 vision transformers and 10 convolutional neural networks to produce a total of 20 ‘votes’. Highly pure (but incomplete) merger samples can be assembled by requiring a unanimous outcome in which 20/20 networks assign a positive merger classification. However, Ferreira et al. (2024a) show that a majority vote ($>$10/20) represents an excellent compromise between purity and completeness, and is hence the threshold adopted in our previous and current works.
@@
Since we ultimately wish to apply the merger classifier to actual survey data obtained by UNIONS, it is important to train mummi with realistic data (Bottrell et al. 2019; Bickley et al. 2024b). Mock $r$-band images of the TNG-selected galaxies are therefore generated with the same image qualities as UNIONS observations (including insertion into real frames to replicate CCD defects and foreground/background objects). With this combined set of information, mummi is able to predict not only the merger status of a galaxy (non-merger, pair or post-merger) but also the time since coalescence ($T_{PM}$) in one of four bins, consolidated for maximal performance: $0.160<T_{PM}<$ 0.16, $0.480.16<T_{PM}<$ 0.48, $0.960.48<T_{PM}<$ 0.96 and $1.760.96<T_{PM}<$ 1.76 Gyr. Note that, in the work presented here, we take $T_{PM}$ = 0 to be the first snapshot after coalescence has occcured, although in practice the actual merger will have happened at some point in the preceding snapshot interval (of $\sim$ 160 Myr duration). After imposing a threshold for prediction accuracy, post-merger galaxies are correctly assigned into these $T_{PM}$ bins 70–80 per cent of the time (Ferreira et al. 2025).
@@ 3.4 The luminosities of merger triggered AGN
All NLAGN in galaxy pairs are successfully matched, typically with $>$100 controls per galaxy. The vast majority ($>$99 percent) of post-mergers are also successfully matched, typically with $>$30 controls per galaxy (the control pool is smaller because it is limited to the UNIONS-SDSS overlap). For the minority of post-merger NLAGN hosts with fewer than five controls the matching tolerances are increased by 0.1 dex in stellar mass and 0.01 in redshift, which yields a successful match for the remaining cases. We then compute a $\Delta L([OIII])$ for each merging NLAGN galaxy, which is the difference (in log space) between the [OIII] luminosity of the merger and the median L([OIII]) of its controls. Thus, the calculation of $\Delta L([OIII])$ is directly analogous to previous calculations of star formation rate enhancements ($\Delta$SFR) in mergers presented in numerous previous works (e.g. Scudder et al. 2012; Bickley et al., 2022; Ferreira et al. 2024b).
@@ 4.2 The AGN-starburst connection
Much of this previous observational work has relied on modelling stellar populations and assessing the epoch at which different physical proceseses (such as starbursts) are occuring. By quantifying the time since merger mummi offers a complementary tool with which to study the sequence of events that are triggered during galaxy-galaxy interactions. In Ferreira et al. (2024b) we showed that the peak of star formation occurs around the time of coalescence. Specifically, SFRs are statistically elevated by a factor of $\sim$ 2 in the closest pairs ($r_{p}<10$ kpc) and the first two post-merger bins ($T_{PM}<0.48$ Gyr). Using reconstructed orbits from the TNG simulation, Patton et al. (2024) calculated that pair separations decrease, on average, by 55 kpc/Gyr. Thus we can conclude that the strongest SFR enhancements begin approximately 0.2 Gyr before coalescence and are in decline by $\sim$ 0.5 Gyr post-merger.
@@ 5 Summary
  • The enhancement in the AGN fraction is long-lived. Different AGN diagnostics are sensitive to accretion events of varied characteristics, for example in terms of luminosity, dust and accretion rate. This is reflected in different quantitative levels of statistical enhancement. However, for both BLAGN and mid-IR selected, a statistical excess is seen out to the longest timescales represented in our sample (1.76 Gyr post-coalescence), see Figure 5. Therefore, compared with the enhancements in SFR, which appear to return to normal levels by $T_{PM}\sim 1$ Gyr (Ferreira et al. 2024b), nuclear accretion experiences a more protracted boost from the merger.
@@
Many of our results are in excellent qualitative agreement with conclusions drawn from cosmological simulations. For example, Byrne-Mamahit et al. (2024) have used the Illustris-TNG simulation to show that the AGN enhancement peaks at coalescence and shows a statistical excess until $T_{PM}\sim$ 1.5 Gyr. As we have also seen in the observations, the excess in AGN in the simulations is also luminosity dependent. Moreover, again using Illustris-TNG, Hani et al. (2020) found a much shorter timescale for SFR enhancement of $\sim$ 500 Myr. Although this is somewhat shorter than the timescale for SFR enhancements that we find in the observations ($\sim$ 1 Gyr; Figure 12 of this paper and Ferreira et al. 2024b) both simulations and observations alike agree qualitatively that post-merger AGN activity is more persistent than enhancements in SFR. However, neither observations nor simulations are able to assign all luminous AGN to mergers (Byrne-Mamahit et al. 2023, 2024). Higher time and spatial resolution in the simulations may be crucial to tackling this outstanding question.

### v2

@@ Abstract
Galaxy mergers have been shown to trigger active galactic nuclei (AGN) in the nearby universe, but the timescale over which this process happens remains unconstrained. The Multi-Model Merger Identifier (mummi) machine vision pipeline has been demonstrated to provide reliable predictions of time post-merger ($T_{PM}$) for galaxies selected from the Ultraviolet Near Infrared and Optical Northern Survey (UNIONS) up to $T_{PM}$=1.76 Gyr after coalescence. By combining the post-mergers identified in UNIONS with pre-coalescence galaxy pairs, we can, for the first time, study the triggering of AGN throughout the merger sequence. AGN are identified using a range of complementary metrics: mid-IR colours, narrow emission lines and broad emission lines, which can be combined to provide insight into the demographics of dust and luminosity of the AGN population. Our main results are: 1) Regardless of the metric used, we find that the peak AGN excess (compared with a matched control sample) occurs immediately after coalescence, at $0.160<T_{PM}<$ 0.16 Gyr. 2) The excess of AGN is observed until long after coalescence; both the mid-IR selected AGN and broad line AGN are more common than in the control sample even in the longest time bin of our sample ($0.96<T_{PM}$ < 1.76 Gyr). 3) The AGN excess is larger for more luminous and bolometrically dominant AGN, and we find that AGN in post-mergers are generally more luminous than secularly triggered events. 4) A deficit of broad line AGN in the pre-merger phase, that evolves into an excess in post-mergers is consistent with evolution of the covering fraction of nuclear obscuring material. Before coalescence, tidally triggered inflows increase the covering fraction of nuclear dust; in the post-merger regime feedback from the AGN clears (at least some of) this material. 5) The statistical peak in the triggering of starbursts occurs contemporaneously with AGN, within 0.16 Gyr of coalescence.
@@ 1 Introduction
For many years, the link between galaxy mergers and the triggering of nuclear accretion was considered highly contentious. Indeed, it had become an almost customary component of any introduction on the subject to point out the slew of papers that both found support for (e.g. Alonso et al. 2007; Woods & Geller 2007; Koss et al. 2010; Ellison et al. 2011; Liu et al. 2012; Sabater et al. 2013; Khabiboulline et al. 2014; Satyapal et al. 2014) and against (e.g. Ellison et al. 2008; Li et al. 2008; Gabor et al. 2009; Darg et al. 2010; Cisternas et al. 2011; Scott & Kaviraj 2014) a link between galaxy mergers and a statistical excess of active galactic nuclei (AGN). Since these pioneering works of over a decade ago, we have learned many lessons, including the importance of large statistical datasets, rigourously matched control samples and multi-wavelength AGN selection. Thus, although the situation at $z>0.5$ is still debated, with some studies supporting a connection between mergers and AGN (e.g. Silverman et al. 2011; Dougherty et al. 2024; La Marca et al. 2024) and others not (e.g. Kocevski et al. 2012; Shah et al. 2020; Silva et al. 2021; Lambrides et al. 2021), modern studies of low redshift galaxies ($z<0.5$) now generally agree that mergers are capable of triggering AGN (e.g. Weston et al. 2017; Goulding et al. 2018; Ellison et al. 2019; Gao et al. 2020; Duplancic et al. 2021; Bickley et al. 2023, 2024a; Li et al. 2023; La Marca et al. 2024).
@@
Despite the broad consensus view that low $z$ mergers can trigger AGN, many details of the merger-AGN connection remain unclear. For example, whether or not merger triggered AGN are more luminous (Koss et al. 2012; Satyapal et al. 2014; Marian et al. 2020; Bickley et al. 2023; Villforth et al. 2017; Villforth 2023; La Marca et al. 2024) or more obscured (Kocevski et al. 2015; Ricci et al. 2017; Satyapal et al. 2017; Koss et al. 2018; Li et al. 2020; Barrows et al. 2023) than secularly triggered AGN, and whether or not feedback is capable of excavating nuclear gas and dust in a so-called blowout phase (e.g. Hou et al. 2020, 2023; Ricci et al. 2021; Bickley et al. 2024a). Definitive answers to these questions remain elusive, with the added complexity that factors such as galaxy mass, star formation rate or the mass ratio of the merger may influence the outcome (e.g. Stemo et al. 2021; Barrows et al. 2023). What seems clear, however, is that the details of the sample selection, both in terms of galaxy (e.g. stellar mass, redshift, orbital geometry and merger mass ratio) and AGN properties (e.g. luminosity and obscuration) all affect the extent to which mergers and AGN are connected, as also found in simulations (e.g. Capelo et al. 2015; Blecha et al. 2018; Steinborn et al. 2018; McAlpine et al. 2020).
@@
In addition to the many galaxy-centric factors (such as gas fraction and morphology) that impact the triggering of AGN during galaxy interactions, another important variable that is expected to influence experimental outcomes is time along the merger sequence. On the one hand, the time evolution of galaxy mergers is a natural axis to investigate with simulations and several works have reported a consensus view that elevated accretion can begin in the encounter phase, but peaks around the time of coalescence (Capelo et al. 2015; McAlpine et al. 2020; Byrne-Mamahit et al. 2023). Most recently, by tracking black hole accretion rates throughout the merger sequence in a large sample drawn from the Illustris-TNG simulation, Byrne-Mamahit et al. (2024) found that the statistical elevation of merger triggered nuclear accretion is long lived, persisting for $\sim$ 2 Gyr post-coalescence.
@@
On the other hand, studying merger properties as a function of time is extremely challenging for observational works. The popular proxy for time in the pre-coalescence regime is the projected separation ($r_{p}$) between pairs of galaxies, which is statistically well motivated by simulations (e.g. Patton et al. 2024). Many observational works have found the greatest enhancement in AGN activity at small pair separations (e.g. Ellison et al. 2011; Khabiboulline et al 2014; Hou et al. 2020; Stemo et al. 2021), motivating studies to push into the post-coalescence regime. Indeed, studies of post-mergers have found that the highest rates of AGN activity occur after coalescence (Ellison et al. 2013; Satyapal et al. 2014; Bickley et al., 2023, 2024a), in agreement with simulations. However, what remains completely unconstrained in the observations is how the frequency of AGN evolves with time after coalescence, both in terms of the epoch at which it peaks, and for how long statistical enhancements persist. The application of artificial intelligence to not only identify mergers (Pearson et al. 2019, 2022; Ferreira et al. 2020; Bickley et al. 2021; Walmsley et al. 2022; Omori et el. 2023) but also predict the time since coalescence of a given post-merger remnant, is just now becoming feasible (Koppula et al. 2021; Pearson et al. 2024).
@@
In order to study the effects of galaxy interactions as a function of time post-merger ($T_{PM}$) we have recently developed the MUlti Model Merger Identifier (mummi), a machine vision based framework that can both find mergers and predict $T_{PM}$ (Ferreira et al. 2024; Ferreira et al. in prep). In our first two papers in the science series, we have used mummi to investigate the post-coalescence time evolution of star formation triggering (Ferreira et al. 2025) and quenching (Ellison et al. 2024). In the work presented here we turn our attention to AGN, in order to assess both the epoch of peak nuclear accretion as well as the duration over which the merger influences the heightened activity.
@@ 2.2 The galaxy pair sample
Our pairs sample is constructed following the methods laid out in Patton et al. (2016), except that we use our new photometrically derived stellar masses that include galaxies with a possible broad line component. The pairs sample is compiled by identifying the closest companion of each galaxy in the SDSS DR7 whose stellar mass is within a factor of 10, whose velocity separation $\Delta v<300$ km s-1 and whose projected separation is $r_{p}<100$ kpc. Furthermore, in order to have a sample that is consistent with the post-mergers (see next sub-section) we require that each galaxy in the pair have a stellar mass of log ($M_{\star}/M_{\odot})\geq 10$. There are 21,235 such galaxies that fulfill all of these criteria and that represent our pairs sample [^2] Strictly, this is a sample of galaxies with close companions, rather than a true sample of pairs, because the closest companions identified do not always find unique pairings. I.e. Galaxy A may have galaxy B as its closest companion, but galaxy B’s closest neighbour may be galaxy C. This explains why we have an odd number of galaxies in the ‘pairs’ sample. Nonetheless, we use the nomenclature ‘pairs’ as it is widely adopted in the literature.. The pairs sample is dominated by major mergers, with two thirds of the sample exhibiting a mass ratio within 2:1.
@@ 2.3 The post-merger sample
The post-merger sample is identified by applying the mummi neural network framework to the Ultraviolet Near Infrared and Optical Northern Survey (UNIONS). Since mummi has been extensively described in Ferreira et al. (2024) and Ferreira et al. (in prep), we give only the briefest of details here.
mummi is a machine vision based merger classifier that uses an ensemble of 10 vision transformers and 10 convolutional neural networks to produce a total of 20 ‘votes’. The classifications use all information in the images including streams, shells and bridges, without explicit (human) guidance on what to look for. Highly pure (but incomplete) merger samples can be assembled by requiring a unanimous outcome in which 20/20 networks assign a positive merger classification. However, Ferreira et al. (2024) show that a majority vote ($>$10/20) represents an excellent compromise between purity and completeness, and is hence the threshold adopted in our previous and current works.
@@
Since we ultimately wish to apply the merger classifier to actual survey data obtained by UNIONS, it is important to train mummi with realistic data (Bottrell et al. 2019; Bickley et al. 2024b). Mock $r$-band images of the TNG-selected galaxies are therefore generated with the same image qualities as UNIONS observations (including insertion into real frames to replicate CCD defects and foreground/background objects). With this combined set of information, mummi is able to predict not only the merger status of a galaxy (non-merger, pair or post-merger) but also the time since coalescence ($T_{PM}$) in one of four bins, consolidated for maximal performance: $0.160<T_{PM}<$ 0.16, $0.480.16<T_{PM}<$ 0.48, $0.960.48<T_{PM}<$ 0.96 and $1.760.96<T_{PM}<$ 1.76 Gyr (see Ferreira et al. in prep). The increasing length of the time bins reflects the increasing difficulty encountered by mummi to make accurate time predictions as the merger progresses.
Note that, in the work presented here, we take $T_{PM}$ = 0 to be the first snapshot after coalescence has occcured, although in practice the actual merger will have happened at some point in the preceding snapshot interval (of $\sim$ 160 Myr duration). After imposing a threshold for prediction accuracy, post-merger galaxies are correctly assigned into these $T_{PM}$ bins 70–80 per cent of the time (Ferreira et al., in prep). This threshold is based on the quality of the probability distributions produced by mummi, and removes any uncertain or spurious classifications mitigating any performance degradation between simulationa and observational domains (Ferreira et al. in prep). Moreover, we found that mummi’s performance was constant across all mass ratios used in the training (from 1:1 down to 1:10).
Once trained, mummi is applied to the UNIONS fifth data release (DR5) of $r$-band imaging. In addition to being the deepest available imaging, using the single band $r$ images means our merger identification is not biased by interaction-induced effects such as starbursts. Using the $r$-band images there ensures both a sensitive selection and one driven purely by morphology.
@@ 3.4 The luminosities of merger triggered AGN
All NLAGN in galaxy pairs are successfully matched, typically with $>$100 controls per galaxy. The vast majority ($>$99 percent) of post-mergers are also successfully matched, typically with $>$30 controls per galaxy (the control pool is smaller because it is limited to the UNIONS-SDSS overlap). For the minority of post-merger NLAGN hosts with fewer than five controls the matching tolerances are increased by 0.1 dex in stellar mass and 0.01 in redshift, which yields a successful match for the remaining cases. We then compute a $\Delta L([OIII])$ for each merging NLAGN galaxy, which is the difference (in log space) between the [OIII] luminosity of the merger and the median L([OIII]) of its controls. Thus, the calculation of $\Delta L([OIII])$ is directly analogous to previous calculations of star formation rate enhancements ($\Delta$SFR) in mergers presented in numerous previous works (e.g. Scudder et al. 2012; Bickley et al., 2022; Ferreira et al. 2025).
@@ 4.2 The AGN-starburst connection
Much of this previous observational work has relied on modelling stellar populations and assessing the epoch at which different physical proceseses (such as starbursts) are occuring. By quantifying the time since merger mummi offers a complementary tool with which to study the sequence of events that are triggered during galaxy-galaxy interactions. In Ferreira et al. (2025) we showed that the peak of star formation occurs around the time of coalescence. Specifically, SFRs are statistically elevated by a factor of $\sim$ 2 in the closest pairs ($r_{p}<10$ kpc) and the first two post-merger bins ($T_{PM}<0.48$ Gyr). Using reconstructed orbits from the TNG simulation, Patton et al. (2024) calculated that pair separations decrease, on average, by 55 kpc/Gyr. Thus we can conclude that the strongest SFR enhancements begin approximately 0.2 Gyr before coalescence and are in decline by $\sim$ 0.5 Gyr post-merger.
@@ 5 Summary
  • The enhancement in the AGN fraction is long-lived. Different AGN diagnostics are sensitive to accretion events of varied characteristics, for example in terms of luminosity, dust and accretion rate. This is reflected in different quantitative levels of statistical enhancement. However, for both BLAGN and mid-IR selected, a statistical excess is seen out to the longest timescales represented in our sample (1.76 Gyr post-coalescence), see Figure 5. Therefore, compared with the enhancements in SFR, which appear to return to normal levels by $T_{PM}\sim 1$ Gyr (Ferreira et al. 2025), nuclear accretion experiences a more protracted boost from the merger.
@@
Many of our results are in excellent qualitative agreement with conclusions drawn from cosmological simulations. For example, Byrne-Mamahit et al. (2024) have used the Illustris-TNG simulation to show that the AGN enhancement peaks at coalescence and shows a statistical excess until $T_{PM}\sim$ 1.5 Gyr. As we have also seen in the observations, the excess in AGN in the simulations is also luminosity dependent. Moreover, again using Illustris-TNG, Hani et al. (2020) found a much shorter timescale for SFR enhancement of $\sim$ 500 Myr. Although this is somewhat shorter than the timescale for SFR enhancements that we find in the observations ($\sim$ 1 Gyr; Figure 12 of this paper and Ferreira et al. 2025) both simulations and observations alike agree qualitatively that post-merger AGN activity is more persistent than enhancements in SFR. However, neither observations nor simulations are able to assign all luminous AGN to mergers (Byrne-Mamahit et al. 2023, 2024). Higher time and spatial resolution in the simulations may be crucial to tackling this outstanding question.